40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Feb 14 17:40:27 2020, anchal, DailyProgress, ISS, Installed ISS on both paths using SR560s IMG_20200214_174127.jpgISS_Effect.pdf
    Reply  Sat Feb 15 18:57:43 2020, rana, DailyProgress, ISS, Installed ISS on both paths using SR560s 
       Reply  Thu Feb 20 17:42:23 2020, rana, DailyProgress, ISS, Installed New ISS on both paths using SR560s ISS_New_Analysis.pdf
          Reply  Sun Feb 23 18:19:32 2020, rana, DailyProgress, ISS, Installed New ISS on both paths using SR560s 
             Reply  Fri Feb 28 14:00:11 2020, anchal, DailyProgress, ISS, Installed New ISS on both paths using SR560s - PROBLEM 
                Reply  Sun Mar 1 16:34:31 2020, rana, DailyProgress, ISS, Installed New ISS on both paths using SR560s - PROBLEM 
                   Reply  Mon Mar 2 16:48:12 2020, anchal, DailyProgress, ISS, Follow-up, porbalem stopped happening, OLTF plot. NISS_OLTF.pdf
Message ID: 2547     Entry time: Mon Mar 2 16:48:12 2020     In reply to: 2546
Author: anchal 
Type: DailyProgress 
Category: ISS 
Subject: Follow-up, porbalem stopped happening, OLTF plot. 

The issue isn't visible right now. I'm not sure what changed but with a similar power level, I'm able to increase the gain on North ISS to much higher than before. Currently, I'm taking RIN measurements (still using in-loop detector) to update the noise budget plot which is taking some time as I'm trying to measure the uncertainty in the measurement as well. So, the step response plot would come later.


OLTF

  • Open-loop transfer function of the North ISS loop was measured by exciting at port B in SR560 and using mode A-B.
  • Measurement was taken as Input A / Output which is equivalent to open loop gain divided by the transfer function of SR560.
  • I measured the transfer function of SR560 separately and hence obtained the OLTF.
  • The first measurement is actually product of Actuator TF, Plant and Detector and I have plotted it as well.
  • This plot shows that a pole of 20 kHz is coming through the plant. It can't be detector or actuator as they have very high bandwidths.
  • At the lower end, we couldn't reach to the point where AC coupling sets in. SR560 manual (at page 12) says it is 30 mHz.
  • Lower UGF is roughly 5 Hz, which the gain margin from lower ugf must be more than 20 dB.
  • Since the power level of laser changes somehow  over day, keeping this OLTF is not really fixed and moves up and down. Hopefully, in the present setting, it doesn't hit oscillations.

 


Witness detector?

The beatnote detector SN101's DC output has electronically railed. This might be due to a disconnection inside which may or may not be intentional. I don't think I should meddle around with that detector so close to the result. I will instead replace the other 1611 detector in the blocked port of beam splitter (11,19) with a thorlabs PDA10CS and use that for RIN measurement. But since, the loops were in modd of being stable today, I decided to go ahead and take the measurement with current settings. This will also work as in-loop measurement for comparison later.


Data

Attachment 1: NISS_OLTF.pdf  44 kB  | Hide | Hide all
NISS_OLTF.pdf NISS_OLTF.pdf
ELOG V3.1.3-